the mean time taken to learn the basics of a software program by all students is 200 minutes with a standard deviation of 20 minutes, the shape of the distribution is unknown. Find the interval that contains the times taken by at least 60% of all students to learn this software program
1
Expert's answer
2020-10-22T18:18:34-0400

By the Chebyshev's inequality: "P(|X-\mu |<k\sigma)>1-\frac{1}{k^2}".

In our case: "1-\frac{1}{k^2}=0.6."

Thus, "k=\sqrt{\frac{1}{0.4}}=1.58."

So, the interval "(\mu-k\sigma, \mu+k\sigma)=(200-1.58*20,200+1.58*20)=("168.4, 231.6)

contains the times taken by at least 60% of all students.